# Horizon-Length Prediction: Advancing Fill-in-the-Middle Capabilities for Code Generation with Lookahead Planning

### v2

@@ Abstract
Fill-in-the-Middle (FIM) has become integral to code language models, enabling generation of missing code given both left and right contexts. However, the current FIM training paradigm, which reorders original training sequences and then performs regular next-token prediction (NTP), often leads to models struggling to generate content that aligns smoothly with the surrounding context. Crucially, while existing works rely on rule-based post-processing to circumvent this weakness, such methods are not practically usable in open-domain code completion tasks as they depend on restrictive, dataset-specific assumptions (e.g., generating the same number of lines as in the ground truth). Moreover, model performance on FIM tasks deteriorates significantly without these unrealistic assumptions.
We hypothesize that NTP alone is insufficient for models to learn effective planning conditioned on the distant right context, a critical factor for successful code infilling. To overcome this, we propose Horizon-Length Prediction (HLP), a novel training objective that teaches models to predict the number of remaining middle tokens (i.e., horizon length) at each step. HLP advances FIM with lookahead planning, enabling models to inherently learn infilling boundaries for arbitrary left and right contexts without relying on dataset-specific post-processing. Our evaluation across different models and sizes shows that HLP significantly improves FIM performance by up to 24% relatively on diverse benchmarks, across file-level and repository-level, and without resorting to unrealistic post-processing methods. Furthermore, the enhanced planning capability gained through HLP boosts model performance on code reasoning. Importantly, HLP only incurs negligible training overhead and no additional inference cost, ensuring its practicality for real-world scenarios.
@@ 1 Introduction
Large Language Models (LLMs) trained on massive source code data have demonstrated significant progress in coding-related tasks (Lozhkov et al., 2024; Guo et al., 2024; DeepSeek-AI et al., 2024; Hui et al., 2024). While natural language generation predominantly follows a Left-to-Right (L2R) approach, Fill-in-the-Middle (FIM), or infilling, is ubiquitous in code completion scenarios. This prevalence stems from the iterative nature of coding, which involves frequent edits and insertions rather than a single left-to-right pass (Bavarian et al., 2022; Fried et al., 2023). In an infilling task, the model is asked to generate the missing code in the middle, conditioned on both the preceding (left) and the succeeding (right) contexts.
The common practice to achieve FIM capability with uni-directional auto-regressive models is to reorder the original sequence of -- into either -- (PSM) or -- (SPM). This reordering allows the middle part to be predicted while conditioned on both left and right contexts as past tokens (Bavarian et al., 2022; Fried et al., 2023). Typically. an <end_of_insertion> (i.e., <eoi>) token is appended to the end to indicate the generation boundary. At training time, the , , and are determined by a random split, and the reordered sequence is fed to the model for standard next-token prediction (NTP).
A key challenge in FIM is to seamlessly connect the generated to the given considering both fluency and semantics, a difficult task for models to learn in practice. As pointed out by the previous work (Bavarian et al., 2022), though model is trained to generate <eoi> when ends and connects to , it often fails to do so at the right place during test time, resulting in generation that does not connect well to . We hypothesize that this challenge stems from the fundamental difference in prediction horizon compared to standard NTP. In NTP, the model only needs to consider a horizon of one token at a time. In contrast, FIM requires the model to plan for a much longer horizon, i.e., the entire length of the missing section. This extended horizon is crucial because the model must generate a sequence that not only follows the left context but also smoothly transitions to the right context, which may be many tokens away. Standard NTP training does not adequately prepare models for this long-horizon planning task. Consequently, models often struggle to maintain coherence over the longer sequences required in FIM, particularly when approaching the transition to the right context. Without effective long-horizon planning, generation frequently falters towards the end of the infill, failing to create a smooth connection with the given .
Most existing FIM benchmarks circumvent the above challenge to some extent by devising rule-based post-processing to truncate redundant parts at the end of the generation (Gong et al., 2024; Zhang et al., 2023; Ding et al., 2023; Wu et al., 2024), as shown in Figure 1. We argue that these post-processing techniques lack practicality as they rely on restrictive and dataset-specific assumptions, such as the exact number of lines of the expected completion (Zhang et al., 2023; Wu et al., 2024) or a specific structure to be met (Ding et al., 2023; Gong et al., 2024). Essentially, we need to develop better models that are capable of spontaneously terminating generation at the correct point with respect to arbitrary left and right contexts.
To this end, we propose Horizon-Length Prediction (HLP) to improve code infilling by teaching models to plan ahead on the number of tokens to be generated (i.e., horizon length). Specifically, given the hidden state of the current token, we introduce an auxiliary training objective to predict the number of future tokens required to complete , in addition to standard next-token prediction (NTP). Unlike rule-based post-processing, HLP is generalizable as it does not require any task-specific knowledge.
Through comprehensive evaluation we demonstrate that HLP achieves up to 24% improvements relatively on diverse FIM benchmarks at both file-level and repository-level, with no access to task-specific post-processing. Moreover, with an emphasis on planning, training with HLP also helps achieve superior model performance on code reasoning. Besides, HLP is also extremely efficient as its training overhead is negligible and it doesn’t have any inference cost.
@@
We highlight that post-processing methods adopted by current benchmarks overestimate existing code LLMs’ FIM performance, and empirically quantify the gap. We further draw attention to models’ long-horizon planning capability as the key to successful code infilling.
We propose Horizon-Length Prediction (HLP), a novel training task that advances fill-in-the-middle capability by teaching LLMs to plan ahead over arbitrarily long horizons. HLP complements the standard next-token prediction by training LLMs to predict the remaining number of future tokens required to complete (i.e., horizon length).
Our evaluation shows that HLP not only improves code infilling by up to 24% across various benchmarks without using any rule-based and/or dataset-specific post-processing, but also enhances performance on code reasoning. HLP is also super efficient as it only incurs negligible training overhead while not adding any inference overhead.
| Post-processing Criteria | |
| RepoEval | Truncate generation to the same number of lines as in ground truth. |
| (Zhang et al., 2023) | |
| CrossCodeLongEval | |
| (Wu et al., 2024) | |
| CrossCodeEval | Truncate generation at the first complete statement. |
| (Ding et al., 2023) | |
| SAFIM | Stop when the target program structure in ground truth is generated. |
| (Gong et al., 2024) |
@@ 2 Post-processing for Fill-in-the-Middle
Most existing FIM works rely on post-processing to truncate code completions generated by LLMs for infilling tasks (Gong et al., 2024; Zhang et al., 2023; Ding et al., 2023; Wu et al., 2024). While such post-processing can enhance the FIM performance, we argue that they are designed based on dataset-specific prior knowledge, and thus fundamentally limited and impractical for real-world scenarios (§2.1). Through evaluation, we show that FIM performance of existing code models drops significantly without post-processing, suggesting that post-processing conceals models’ inability to determine the end of insertion (§2.2). Furthermore, we show that plausible generation of without careful planning can easily lead to the failure in connecting to at the end, which can not be mitigated even with post-processing (§2.3). We believe that post-processing leads to an overestimation of infilling capability of existing code LLMs and more generalizable techniques are in need to advance LLMs’ FIM performance.
2.1 Post-processing Requires Task-specific Knowledge
Post-processing methods adopted by recent FIM benchmarks typically assume a certain completion type and perform rule-based truncation accordingly. Table 1 summarizes the post-processing criteria of four popular FIM benchmarks, highlighting the specific rule used for each dataset. These criteria do not transfer across datasets, nor are they generalizable to FIM in the real-world scenario where both left and right contexts can be arbitrary. Given the complexity of programming languages, it is infeasible to devise any rule-based post-processing for general code infilling in open-domain scenarios. A more practical and effective solution is to let the model itself decide when to stop by learning from the massive data.
2.2 Post-processing Conceals LLMs’ Inability of Connecting to Suffix
To further demonstrate to what extent post-processing conceals LLMs’ inability of connecting to , we conduct a comprehensive experiment on SAFIM. We compare FIM performance of four different code LLMs, with or without post-processing. As shown in Table 2, after removing post-processing, we see pass@1 drops by up to 13.8% across all models, revealing that post-processing seriously obfuscates LLMs’ inherent inability to connect seamlessly to during infilling.
2.3 Fill-in-the-Middle Requires Planning Capability of LLMs
We argue that FIM requires planning from LLMs for deeper reasons beyond simply predicting the <eoi> token. Figure 2 provides an illustrative example that highlights the importance of planning ahead in FIM tasks. In this example, compared with the ground truth (i.e., in Reference), the model generation (i.e., in Answer) does not correctly connect to . To elaborate, since the beginning of is a function call (Recognizer()) that can only be accessed through a specific object (speech), the model has to end the generation with speech. to properly connect to , but it fails to do so in Answer.
Interestingly, the model has demonstrated its knowledge of how to call the function by generating speech.Recognizer() in Answer. However, without careful planning, it prematurely writes this call before other necessary code (the assignment statement of self.voice). This example illustrates that the real challenge lies in planning the entire generation, rather than understanding individual components. Furthermore, this case demonstrates that post-processing, despite taking advantage of task-specific prior knowledge, cannot adequately address such planning failures. Truncating the code after speech. in Answer would result in loss of other necessary code for correctness. This observation underscores the irreplaceable importance of planning ahead in code infilling.
@@ 3 Horizon-Length Prediction
Given a document that contains tokens , existing FIM training scheme can be divided into three steps: (1) Split the document into three parts: -- 111We opt to use PSM setting in this work given our base models (DeepSeek-Coder (Guo et al., 2024) and StarCoder2 (Lozhkov et al., 2024)) were both pre-trained with PSM setting. However, we expect that our method is generalizable to SPM setting as well., (2) Construct a new FIM-style document by reordering the three parts as --, and (3) Conduct next-token prediction (NTP) training on the document .
@@
While NTP has provided LLMs with necessary supervision signals to learn Fill-in-the-Middle (i.e., starting generation to match and stopping at the right place to connect to ), it is shown in our analysis (§2) that LLMs trained with NTP alone do not learn this task well in practice. This is likely due to models trained with NTP only learns to predict with a horizon length of only 1 (i.e., next token), and thus infilling which typically requires planning over a much longer horizon length could not be learnt effectively.
Horizon-Length Prediction (HLP). To mitigate this issue, we propose adding an auxiliary training objective, namely horizon-length prediction (HLP), to improve the planning capabilities of LLMs over long horizons. Specifically, given the hidden state of current token, the model is tasked by HLP to predict the number of future tokens required to complete , as shown in Figure 3.
In detail, the FIM-style document defined in Eq. (1) contains three different parts, including , , and , and HLP is applied to part as it is the major focus of FIM tasks. Recall that has tokens, ranging from to . At each position in , HLP aims to predict the number of future tokens required to complete the , which is . Considering that the context window size of LLMs can be infinite in theory and so is , instead of formatting HLP as a classification task with a fixed set of discrete values as labels, we make it a regression problem by taking the normalized count of future tokens as the target:
As such, the target is always within the interval regardless of the model’s context window size.
@@
where is the tunable weight. In experiments, we set , which achieves good results across benchmarks empirically.
Why should HLP work? Theoretically, next-token prediction (NTP) is myopic in the sense that it does not take into account future constraints when predicting the immediate next token, as it only optimizes for local token-by-token decisions. While this is less of a problem for left-to-right code generation, in FIM there is always such a constraint that the middle part must connect smoothly to the right context. Standard NTP only tackles this constraint at the very last <eoi> token. However, at test time, the <eoi> token can sometimes be completely missing due to accumulated errors in preceding tokens which are not trained with any end-of-insertion information (Gong et al., 2024). In contrast, with HLP, every token gets a sense of how far away it is from connecting to the right context (i.e., global sequence-level planning), leading to a more reliable closing of the infilling generation. In other words, while NTP only provides a sparse signal from <eoi> token prediction at the end of generation, HLP enables a consistent optimization by providing an auxiliary learning signal at every generation step.
Overhead Analysis. While HLP introduces the additional during training, the number of added parameters is < 0.01% of the base model, which incurs almost zero training overhead. Furthermore, the additional head can be discarded during inference, leading to no inference overhead.
@@ 4 Experiments
Training. We conduct continual pre-training on a set of code LLMs of different model families and sizes to validate the effectiveness of HLP. Specifically, DeepSeek-Coder-Base 1.3B/6.7B (Guo et al., 2024) and StarCoder2 3B/7B (Lozhkov et al., 2024) are involved in our experiments. We use AdamW (Loshchilov & Hutter, 2019) as the optimizer with and . We use a cosine learning rate scheduler with a peak learning rate equal to that at pre-training end. All models are trained for 200K steps with a warm-up period over the first 3,000 steps. The global batch size is 512.
Dataset. We use a subset of the-stack-v2-train-smol (Lozhkov et al., 2024) for continual pre-training which includes Python, Java, C++, and C#. In line with existing works (Guo et al., 2024; Lozhkov et al., 2024), FIM rate is set to . We employ Best-fit Packing (Ding et al., 2024) to group multiple files into each training sequence while masking out cross-file attention. The -- split is applied to each file independently rather than the whole training sequence.
@@
| Algorithmic | Algorithmic | Control | API | ||
@@
NTP + HLP: our newly proposed pre-training scheme that incorporates horizon-length prediction (HLP) objective with next-token prediction (NTP) objective.
Throughout this section, we determine the end of generation solely based on <eoi> predicted by the model during evaluation, without any rule-based post-processing, unless otherwise specified in §4.4. We also conduct statistical analysis to validate our findings. Specifically, we conduct paired t-tests between the baseline (NTP) and our approach (NTP +HLP) across all benchmarks. We denote results that are statistically significant () by a superscript *.
4.1 Syntax-Aware and Multilingual Code Fill-in-the-Middle
We use SAFIM (Gong et al., 2024), a syntax-aware and multilingual code Fill-in-the-Middle benchmark, to evaluate the effectiveness of HLP. SAFIM focuses on syntax-aware completions of program structures, covering algorithmic block (i.e., Algo and Algo), control-flow expression (i.e., Control), and API function call (i.e., API). It consists of 17,720 examples from four different programming languages, including Python, Java, C++, and C#. SAFIM employs execution-based evaluation and reports pass@1 as the evaluation metric. As shown in Table 3, compared with NTP only, adding HLP achieves up to 5% improvements on average across all the studied code LLMs. Specifically, HLP consistently improves LLMs’ performance on completions of various program structures. Furthermore, since SAFIM is a multilingual benchmark, our evaluation also shows that incorporating HLP can improve FIM performance across different languages.
4.2 Repository-Level Cross-File Code Fill-in-the-Middle

### v3

# Planning-Aware Code Infilling via Horizon-Length Prediction

@@ Abstract
Fill-in-the-Middle (FIM), or infilling, has become integral to code language models, enabling generation of missing code given both left and right contexts. However, the current FIM training paradigm which performs next-token prediction (NTP) over reordered sequence often leads to models struggling to generate content that aligns well with the surrounding context. We hypothesize that NTP alone is insufficient for models to learn effective planning conditioned on the distant right context, a critical factor for successful code infilling. To overcome this, we propose Horizon-Length Prediction (HLP), a novel training objective that teaches models to predict the number of remaining middle tokens at each step. HLP advances FIM with lookahead planning, enabling models to inherently learn infilling boundaries for arbitrary left and right contexts without relying on dataset-specific post-processing. Our evaluation across different model families and sizes shows that HLP significantly improves FIM performance by up to 24% relatively on diverse benchmarks, across file-level and repository-level. Furthermore, the enhanced planning capability gained through HLP boosts model performance on code reasoning. Importantly, HLP incurs negligible training overhead and no additional inference cost, ensuring its practicality for real-world scenarios.
Planning-Aware Code Infilling via Horizon-Length Prediction
Yifeng Ding1††thanks: Work done during a research internship at AWS AI Labs. Hantian Ding2 Shiqi Wang2 Qing Sun2 Varun Kumar2 Zijian Wang2 University of Illinois Urbana-Champaign1 AWS AI Labs2 yifeng6@illinois.edu {dhantian,zijwan}@amazon.com
@@ 1 Introduction
Fill-in-the-Middle (FIM), or infilling, has become essential for modern code development, where programmers frequently need to insert or modify code between existing sections rather than writing linearly from start to end (Bavarian et al., 2022; Fried et al., 2023). While large language models have shown remarkable capabilities in code generation, the FIM task poses unique challenges that go beyond traditional left-to-right generation. A model must not only generate code that follows from the preceding context (), but also smoothly connect to the subsequent code () – a task that requires careful planning and foresight.
Current approaches to FIM typically reorder the input sequence and apply standard next-token prediction (NTP) training (Lozhkov et al., 2024; Guo et al., 2024; DeepSeek-AI et al., 2024; Hui et al., 2024). However, as illustrated in Figure 1, this methodology has a fundamental limitation: models struggle to maintain coherence over longer sequences and often fail to create smooth transitions to the right context (Bavarian et al., 2022). While existing benchmarks attempt to address this through rule-based post-processing (e.g., truncating generated code based on line counts (Zhang et al., 2023; Wu et al., 2024) or program structure (Ding et al., 2023; Gong et al., 2024)), such methods rely on dataset-specific assumptions that do not generalize to real-world scenarios where both left and right contexts can be arbitrary.
| Post-processing Criteria | |
| RepoEval (Zhang et al., 2023) | Truncate generation to the same number of lines as in ground truth. |
| CrossCodeLongEval (Wu et al., 2024) | |
| CrossCodeEval (Ding et al., 2023) | Truncate generation at the first complete statement. |
| SAFIM (Gong et al., 2024) | Stop when the target program structure in ground truth is generated. |
We hypothesize that the core challenge lies in the model’s limited ability to plan ahead while filling in the middle. Standard next-token prediction training only requires considering one token at a time, but successful code infilling demands awareness of the entire missing section to ensure both local coherence and proper connection to the right context. This planning capability is particularly crucial because generated code must satisfy both syntactic requirements and semantic constraints from both surrounding contexts.
To address this limitation, we propose Horizon-Length Prediction (HLP), a novel training objective that teaches models to predict the number of remaining tokens needed to complete the middle section at each generation step. HLP advances NTP by encouraging models to develop awareness of the generation horizon and plan accordingly. Unlike post-processing approaches, HLP is generalizable as it does not rely on any task-specific knowledge. Instead, it strengthens models’ innate ability to plan and execute coherent code completions.
Our comprehensive evaluation across different models and model sizes demonstrates that HLP significantly improves FIM performance, achieving up to 24% relative improvement on diverse benchmarks at both file-level and repository-level without relying on any dataset-specific post-processing. Furthermore, the enhanced planning capability gained through HLP training also boosts model performance on code reasoning tasks. Importantly, HLP incurs negligible training overhead and no additional inference cost, making it practical for real-world deployment.
@@
We identify planning capability as a fundamental bottleneck in current FIM approaches and quantitatively demonstrate how post-processing methods in existing benchmarks obscure this critical limitation.
We propose Horizon-Length Prediction (HLP), a novel training objective that advances fill-in-the-middle capability by teaching models to plan over arbitrarily long horizons, with negligible training and zero inference overhead.
We demonstrate up to 24% improvement in FIM performance across multiple benchmarks and model families without relying on any post-processing.
We show that HLP’s benefits extend beyond FIM to improve performance on code reasoning tasks, and present various analyses that illuminate the underlying mechanism that enable these improvements.
@@ 2 Post-processing for Fill-in-the-Middle
Most existing FIM works rely on post-processing to truncate code completions generated by LLMs for infilling tasks (Gong et al., 2024; Zhang et al., 2023; Ding et al., 2023; Wu et al., 2024). While such post-processing can enhance the FIM performance, we argue that they fundamentally depend on dataset-specific assumptions that make them impractical for real-world scenarios (§2.1). Through evaluation, we show that the performance of FIM existing code models drops significantly without post-processing, suggesting that post-processing conceals the fundamental struggles of models with code filling (§2.2). Furthermore, we show that this limitation stems from models’ inability to plan coherent completions given a fixed – a challenge that post-processing alone cannot address.
2.1 Post-processing Requires Task-Specific Knowledge
Post-processing methods adopted by recent FIM benchmarks typically assume a certain completion type and perform rule-based truncation accordingly. Table 1 summarizes the post-processing criteria of four popular FIM benchmarks, highlighting the specific rule used for each dataset. These criteria do not transfer across datasets, nor are they generalizable to FIM in the real-world scenario where both left and right contexts can be arbitrary. Given the complexity of programming languages, developing universally applicable post-processing rules for infilling is infeasible. Instead, models need to learn the intrinsic patterns that make for good completions.
@@ 2.1 Post-processing Requires Task-Specific Knowledge
2.2 LLMs Fail to Plan Coherent Completions
To further demonstrate to what extent post-processing conceals LLMs’ inability of connecting to , we conduct a comprehensive experiment on SAFIM. We compare FIM performance of four different code LLMs, with or without post-processing. As shown in Table 2, removing post-processing leads to up to a substantial 13.8% Pass@1 drop across all models. This reveals that current models have not truly mastered the fundamental task of generating code that properly connects and contexts. Post-processing creates an illusion of competence by artificially “fixing” problematic generations rather than addressing the core limitation.
2.3 FIM Requires Planning Capability
The challenges in FIM stem from models’ inability to plan coherent completions. Consider the example in Figure 1, where a model needs to implement a speech recognition initialization: while the model demonstrates knowledge of the required components (using Recognizer), it fails to order them properly. Without careful planning, it prematurely places the Recognizer call, leading to both syntactic and semantic errors. Importantly, this type of failure cannot be fixed through post-processing, as truncating the generation would lose essential statements while keeping it results in invalid code.
Such an example illustrates that successful FIM requires not only knowledge of individual components but also the model’s ability to plan coherent sequences that smoothly connect to both contexts. The model must reason about the entire completion considering both local coherence and global structure, and that points to a clear need for models to develop intrinsic planning capabilities to succeed at FIM.
@@ 3 Horizon-Length Prediction
Given a document that contains tokens , existing FIM training scheme can be divided into three steps: (1) Split the document into three parts: -- 111We opt to use PSM setting in this work given our base models DeepSeek-Coder and StarCoder2 were both pre-trained with PSM. We expect that our method is generalizable to SPM setting as well., (2) Construct a new FIM-style document by reordering the three parts as --, and (3) Conduct next-token prediction (NTP) training on the document .
@@
While NTP provides basic FIM capabilities, it has a fundamental limitation: the model only learns to predict one token at a time without developing awareness of the overall horizon. This makes it difficult to plan coherent sequences that properly connect to both contexts.
Horizon-Length Prediction (HLP). To mitigate this issue, we propose horizon-length prediction (HLP) as an auxiliary training objective. As shown in Figure 2, the key idea is to teach models to predict the number of remaining tokens needed to complete the section at each generation step. This creates an explicit training signal for planning awareness.
Specifically, at each position in , HLP predicts
where is the total length of . This normalized value represents the portion of that remains to be generated and ensures the target is always within interval regardless of the model’s context window size.
@@
where is the tunable weight. In experiments, we set , which achieves good results across benchmarks empirically. We further study the relationship between the values of and the FIM performance of models trained with HLP. As discussed in Appendix A.1, the performance of HLP is robust to different values of .
Overhead Analysis. While HLP introduces the additional during training, the number of added parameters is < 0.01% of the base model, which incurs almost zero training overhead. Furthermore, the additional head will not be used during inference, leading to zero inference overhead.
@@ 4 Experiments
| Algo | Algo | Control | API | ||
@@
Training. We conduct continual pre-training on a set of code LLMs of different model families and sizes to validate the effectiveness of HLP. Specifically, DeepSeek-Coder-Base 1.3B/6.7B (Guo et al., 2024) and StarCoder2 3B/7B (Lozhkov et al., 2024) are involved in our experiments. We use AdamW (Loshchilov and Hutter, 2019) as the optimizer with and . We use a cosine learning rate scheduler with a peak learning rate equal to that at the pre-training end. All models are trained for 200K steps with a global batch size of 512.
Dataset. We use a subset of the-stack-v2 (Lozhkov et al., 2024) including Python, Java, C++, and C#. Following existing works (Guo et al., 2024; Lozhkov et al., 2024), FIM rate is set to . We employ Best-fit Packing (Ding et al., 2024) to group multiple files into each training sequence while masking out cross-file attention. The -- split is applied to each file independently rather than the whole training sequence.
@@
NTP + HLP: our newly proposed objective that incorporates horizon-length prediction (HLP) objective with next-token prediction (NTP) objective.
Throughout this section, we determine the end of generation solely based on <eoi> generated by the model without any rule-based post-processing, unless otherwise specified (§4.4). We also conduct statistical analysis: results marked with * are statistically significant () based on paired t-tests.
4.1 Syntax-Aware Multilingual Code FIM
We use SAFIM (Gong et al., 2024), a syntax-aware and multilingual code Fill-in-the-Middle benchmark, to evaluate the effectiveness of HLP. SAFIM focuses on syntax-aware completions of program structures, covering algorithmic block (i.e., Algo and Algo), control-flow expression (i.e., Control), and API function call (i.e., API). It consists of 17,720 examples from four different programming languages, including Python, Java, C++, and C#. SAFIM employs execution-based evaluation and reports pass@1 as the evaluation metric. As shown in Table 3, compared with NTP only, adding HLP achieves up to 5% improvements on average across all the studied code LLMs. Importantly, the improvement is consistent across languages and program structures, demonstrating HLP’s ability to enhance FIM capabilities regardless of language or completion context.
4.2 Repository-Level Cross-File Code FIM